Please help me with this linear equation: 7(2x-3)-4(x+5)=8(x-1)+3 I do all of my work on computers that is why i don't have a textbook . This book is called Algebra DeMYSTiFieD a self teaching guide! Thank you!
Answer by stanbon(75887)

7(2x-3)-4(x+5)=8(x-1)+3
Use the distributive property to get:
14x-21-4x-20=8x-8+3
Combine like terms to get:
10x-41=8x-5
Subtract 8x and add 41 to both sides of the equation to get:
2x=36
Divide both sides by 2 to get:
x=18
-------------------
Cheers,
Stan H.